Assume the following:

• Required reserve ratio of 0.15

• New deposit of $2500.00

How much deposit creation will be created from this new deposit?
Answer #1

Money multiplier=1/required reserve ration=1/0.15=6.67

Deposit creation=new deposit×money multiplier

=$2500.00×6.67=$16,675

From new deposit =16,675-2500=14,175(money created)